OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

-3 = 7(1)+b -3 = 7 + b subtract 7 from both sides -10 =b So the answer is A

OpenStudy (anonymous):

Did that help you or do you need me to explain what i did?

OpenStudy (anonymous):

I hope

OpenStudy (anonymous):

Do you understand what you plug in?